# Getting Explicit Instruction Right

### v1

@@ Abstract
There has been substantial public debate about the potentially deleterious effects of the long-run move to “inquiry-based learning” in which students are placed at the center of an educational journey and arrive at their own understanding of what is being taught. There have been numerous calls for a return to “direct” or “explicit” instruction. This paper focuses on identifying the causal effect of correctly implementing explicit instruction on student performance in standardized tests. We utilise a unique setting in Australia—a country in which all students in grades 3, 5, 7, and 9 undergo annual basic skills tests (“NAPLAN”). We use a synthetic control approach to study the effect of introducing Explicit Instruction in Charlestown South Public School—a median-performing school—on Year-3 and Year-5 NAPLAN scores in Reading and Numeracy. Importantly, this is achieved via peer modelling, with Charlestown teaching staff sitting-in during the classes of a high-performing explicit-instruction school. We find that the performance gains are substantial and persistent.
@@ 1 Introduction
While the source of this decline is hard to pin down, many have speculated that a long-run trend in the way students are taught is to blame. This trend has involved a move away from teachers directly imparting learning to students, toward so-called “inquiry-based learning” in which students are placed at the center of an “educational journey” where they arrive at their own understanding of what is being taught. This approach emerged from the “constructivist” movement in education which emphasizes hands-on learning aimed at solving “real-world” problems. This has led to numerous calls for a return to “direct” or “explicit” instruction (Hammond, 2019). This is sometimes called a “back-to-basics” approach to education.
@@
The purpose of this paper is not to litigate the cause of declining educational performance in many advanced economies. Rather, we focus on identifying the causal effect of correctly implementing the explicit instruction pedagogy on student performance in standardized tests. To do this, we utilize a unique setting in Australia—a country in which all students in grades 3, 5, 7, and 9 undergo annual basic skills tests (“NAPLAN”). We find strong causal evidence that a correct implementation of explicit instruction improves student performance and the magnitude is substantial, with point estimates ranging between 0.9 and 2.7 standard deviations. We emphasise the implementation aspect and build the counterfactual outcomes using a pool of schools that includes schools already integrating elements of explicit instruction in their pedagogical approach.
Specifically, we use a synthetic control approach to study the effect of introducing Explicit Instruction in a median-performing school—Charlestown South Public School, New South Wales—on the school’s Year-3 and Year-5 NAPLAN scores, both in Reading and Numeracy (see Table 5 for more information on the performance percentiles of Charlestown South and the key schools used to construct synthetic Charlestown South). Each year, all Australian students in grades 3, 5, 7, and 9 sit a basic skills tests known as “NAPLAN” (the National Assessment Program Literacy and Numeracy). These tests are administered each May in all schools across the country, and are intended to provide a snapshot of students’ current abilities in reading, writing, language (spelling, grammar and punctuation), and numeracy.
@@
There is a large and multi-disciplinary literature providing evidence of positive sign on the effectiveness of explicit instruction and similar pedagogical approaches. While Explicit Instruction is defined in various ways throughout the literature, we follow the definition from Archer and Hughes (2010)—a classic textbook on this topic. See Section 2.1 for details. Relevant studies in the education literature include Andreassen and Bråten (2011), on reading for primary school students, Doabler et al. (2015), on mathematics in Kindergarten, and Kroesbergen and Luit (2003) and Powell et al. (2021), on using explicit instruction to teach mathematics to primary and secondary school students with special needs, respectively. For reviews focusing on the education literature, see Archer and Hughes (2010), Hughes et al. (2017), and Stockard et al. (2018)’s comprehensive meta-analysis.
@@
Economists have also been interested in assessing the impact of explicit and structured teaching approaches on student outcomes. Difference‐in‐differences studies of the “Literacy Hour” and synthetic-phonics reforms find substantial gains in early reading outcomes (approx. $0.2-0.3$ standard deviations; s.d.) and long-run benefits for low-performing students (Machin and McNally, 2008; Machin et al., 2018). Using field experiments in the US, Fryer (2014) provides evidence that importing the direct instruction practices of charter schools—extended instructional time, data‐driven teaching, intensive tutoring—into public schools raises mathematics achievement by up to 0.18 s.d. one year after the implementation. Schwerdt and Wuppermann (2011) estimates a fixed-effects model on US data and finds that approaches where the teacher takes centre stage and presents the content in a structured format outperform problem‐solving approaches. Jacob and Lefgren (2004) uses a regression discontinuity design and focusing on mathematics, and Banerjee et al. (2007) and Piper et al. (2018) running RCTs in India and Kenya, respectively, to study the effect of structured approaches on numeracy and literacy skills—all finding positive impacts. Other examples of studies related to structured-teaching interventions are Romero and Sandefur (2022), an RCT where 93 Liberian schools outsourced were randomly-assigned to be managed by private operators—who used scripted, highly-structured curricula—and Duflo et al. (2024), where where one of the treatment arms of their RCT in Ghana was based on remedial tutoring that uses structured lesson guides. They also find positive impacts.
@@
Our paper adds to this literature in several ways. First, it shows the importance of a correct and high-fidelity implementation of Archer and Hughes (2010)’s Explicit Instruction approach. In our case study, this is achieved by the teachers of the treated school emulating the teachers of a successful Explicit-Instruction school. Specifically, the principal and the teachers of the treated school visited the successful school for three days and learned how to best implement Explicit Instruction by sitting-in during their classes (called “modelling”, in the education literature). This seems not only to be an effective solution, but also a cost-effective one, compared—for instance—to hiring dedicated coaching staff for long periods of time (see Section 2.3), and particularly in high-income countries like Australia, where the cost of labour is high. The fact that this is happening outside of an experimental setting—and within a public-school context—should also reassure policy-makers that the performance gains we estimate are attainable.
@@
Second, our intervention of interest does not target a specific subject (e.g., Machin and McNally, 2008; Doabler et al., 2015; Machin et al., 2018) or student sub-group (e.g., Torgesen et al., 2010; Powell et al., 2021; Cortes et al., 2015), instead changing the school pedagogy across all teaching areas and benefiting all students. This allows our evidence to provide insight on the broad benefits of a school-wide adoption of Explicit Instruction, compared to a partial or targeted one. Studies of charter schools and their practices (e.g., Dobbie and Fryer, 2013; Fryer, 2014) are similar to our study in this respect, but do not focus strictly on features of Explicit Instruction.
@@
Third, we are able to follow students for 8 years after the implementation, while even most studies with longer follow-up times are limited to less than two years (e.g., Jacob and Lefgren, 2004; Banerjee et al., 2007; Torgesen et al., 2010). This allows us to (i) learn that transitioning from one pedagogical approach to another can take time, and (ii) that that the benefits of Explicit Instruction are persistent in the long run. Machin and McNally (2008) follow students between 1996 and 1998, but focus exclusively on impacts of reading and English skills.
@@
The paper proceeds as follows. Section 2 provides background on Explicit Instruction and the Charlestown South Case Study. Section 3 discusses our data, while section 4 outlines our (synthetic control) approach to causal inference. Section 5 presents our results, and section 6 concludes. Various robustness checks are contained in the appendix.
@@ 2.1 The Pillars of Explicit Instruction
While Archer and Hughes (2010) define sixteen Explicit Instruction principles, we report the five essential components identified in Hughes et al. (2017). The first essential component consists in segmenting complex skills, where complex tasks are divided into smaller simpler units. The second is to indicate important features of the content by either showing, if an action is being taught, or telling, i.e. thinking aloud, if a concept is being taught.
@@
Note that, while some components of Explicit Instruction are found in other pedagogical approaches, Explicit Instruction is distinct from similar-sounding pedagogies such as “Direct Instruction” and “Direct Explicit Instruction” (Hughes et al., 2017). For instance, Direct Instruction includes scripted lessons—which are absent in Explicit Instruction. Moreover, Direct Explicit Instruction and Direct Instruction include both curricular (what to teach) and instructional (how to teach it) elements, while Explicit Instruction purely focuses on instruction (Hughes et al., 2017).
@@ 2.3 Other attempts at introducing Explicit Instruction in NSW
Before Charlestown South started implementing Explicit Instruction, the New South Wales public-school system had already introduced several structured-teaching programmes that feature al least some principles of Explicit Instruction. We provide some examples below.
The Early Action for Success (EaFS) policy was launched in 2012 across the Kindergarten, Year 1 and Year 2 classrooms of 228 low-performing schools. It provides target schools with “instructional leaders”, specially-hired teachers whose role is to train teacher staff in implementing explicit instruction routines both targeting the whole classroom and small groups of students who are not performing well (Centre for Education Statistics and Evaluation, 2016). The Language, Learning and Literacy programme (L3) was piloted in 2010 and scaled to more than 900 Kindergarten classrooms by 2014, providing teachers with scripted, thirty-minute sequences on guided reading and writing (Centre for Education Statistics and Evaluation, 2015). The Focus on Reading was rolled out state-wide in 2011 and trained teachers Years 3–6 teachers turn comprehension lessons into a deliberate, step-by-step routine. (NSW Department of Education and Communities, 2013). All these initiatives are in line with research on evidence-based best practices in teaching conducted by the NSW Department of Education’s Centre for Education Statistics and Evaluation (CESE; New South Wales. Centre for Education Statistics and Evaluation, 2013), which recommended explicit/direct teaching approaches based on a review of the literature and was the basis of the change in pedagogical approach in NSW public schools marked by NSW Government (2013).
@@ 3 Data
The NSW Department of Education dataset gives us information on: each school’s geo-location (both encoded as latitutde-longitude coordinates and as postcodes), whether it is academically selective, how remote is its surrounding area, whether it includes at least one Opportunity Class (a form of “tracking”), the year in which it hired the first teacher, school-level data on student attendance, the school’s type (primary or secondary) and its sub-type (e.g. Kindergarten to Year-6), whether it is single-sex or coeducational, if it is open until late in the and if it has a healthy canteen. It also reports time-varying school-level information such as the Index of Community Socio-Educational Advantage (ICSEA) 111 The ICSEA is an index that summarizes information about the occupation and earnings of children’s parents within a given school, how remote the school is and the percentage of indigenous student enrolled. , the attendance rate, the share of girls, the share of indigenous students, the share of students with a language background other than English, the average teacher-student ratio.
4 The Synthetic Control Method
The Synthetic Control Method (SCM) is a method first introduced by Abadie and Gardeazabal (2003) to study the effect of conflict on economic output, using episodes of terrorism in the Basque Country in the 1960s as a case study. They developed this method to add a quantitative dimension to case study analysis, which was previously purely qualitative. In this section, we introduce the theory of synthetic control as applied our own case study, i.e. the introduction of Explicit Instruction at Charlestown South Public School in 2014 and how it affected NAPLAN scores.
First, we need to select a pool of schools that we judge as being similar to Charletown South, but were not subject to the same intervention—here, learning the Explicit Instruction pedagogy from a successful school. Among all primary public NSW schools, we selected schools that have some key features that are identical to Charlestown South. We exact-match these features as we see them as too important to be matched using the minimum-distance approach embedded in SCM (see the next paragraph). They have the same level of remoteness (ABS, July 2016), “Major Cities of Australia”; they are co-educational; they cover Kindergarten to Year 6, and their student body comprises between 10-15% of students with a language background other than English (Charlestown South’s is 12%). We are left with 108 schools which, in the language of synthetic control, are referred to as donor schools.
Fourth, we construct a synthetic control. This involves creating a “synthetic” version of Charlestown South as similar as possible as the real Charlestown South. Since we cannot observe Charlestown South had it not been treated in 2014, we use its synthetic version to approximate this counterfactual scenario. This allows us to estimate the treatment effect of correctly implementing Explicit Instruction over time by comparing Charlestown South after treatment with its synthetic version. The synthetic control is built by finding a “mix” (the convex hull) of donor schools such that associated donor school weights (i) leading to pre-treatment paths of NAPLAN scores “most similar” to Charlestown South’s and (ii) produce school features of synthetic Charlestown South that are “most similar” to those of the real one.
4.1 Estimation
Formally, we observe the outcome of interest $Y_{it}$—the NAPLAN scores—for each school $i$ and year $t$. We also observe $K$ covariates (or features) per each school, $X_{1i},\ldots,X_{Ki}$, which are themselves unaffected by the intervention. These covariates are selected for their capacity to affect NAPLAN scores. If a covariate changes over time for a given school, rather than being a fixed feature, we take the within-unit mean for that covariate—following Abadie and Gardeazabal (2003).
In the matrix notation used below: (i) the $K\times 1$ vectors $\mathbf{X}_{1},\ldots,\mathbf{X}_{N}$ contains the values of the covariates of units $i=1,\ldots,N$, respectively, and (ii) the $K\times(N-1)$ matrix, $\mathbf{X}_{0}=\left[\mathbf{X}_{1}\cdots\mathbf{X}_{N-1}\right]$, collects the values of the covariates of the $N-1$ untreated (or donor) schools.
We follow (Abadie and Gardeazabal, 2003) and Abadie et al. (2010) and choose weights $\mathbf{W}^{*}=\left(w_{1}^{*},\ldots,w_{N-1}^{*}\right)^{\prime}$ that minimize the Euclidean distance between the covariates of treated and untreated schools in the pre-treatment periods, while restricting the SC weights to being positive and summing up to one.
Formally,
$\min_{\bm{W}}\left\|\mathbf{X}_{1}-\mathbf{X}_{0}\bm{W}\right\|$ $=\sqrt{\left(\mathbf{X}_{1}-\mathbf{X}_{0}\mathbf{W}\right)^{\prime}\mathbf{V}\left(\mathbf{X}_{1}-\mathbf{X}_{0}\mathbf{W}\right)}$ (1)
$=\left(\sum_{k=1}^{K}v_{k}\left(X_{kN}-w_{1}X_{k1}-\cdots\right.\right.\left.\left.-w_{N-1}X_{kN-1}\right)^{2}\right)^{1/2}$ (2)
$\text{where }\mathcal{W}=\left\{\left(w_{1},\ldots,w_{N-1}\right)^{\prime}\right.\text{subject to }$ $\hskip 5.0ptw_{1}+\cdots+w_{N-1}=1,$ (4)
$\hskip 5.0ptw_{i}\geq 0,$ (5)
$\text{with }i=1,\ldots,N-1.\}$ (6)
$\mathbf{V}$ is a $(K\times K)$ symmetric and positive semidefinite matrix. It is another matrix of weights. Indeed, one the one hand, $\bm{W}$ regulates how much importance each donor school has in the newly created synthetic Charlestown South—what share of synthetic Charlestown South was donated by each donor school (notice the $i$ index in $w_{i}$, which stands for school).
On the other hand, $\mathbf{V}$ sets the importance that each covariate has in constructing synthetic Charlestown South. To this end, $\mathbf{V}=\{v_{1}+\cdots+v_{K}\}$ is chosen so that it best predicts $Y_{Nt}(0)$, in the sense that it minimises the Mean Squared Prediction Error (MSPE) between the pre-treatment outcome of the treated school over time ($Y_{Nt}$, $t\leq T_{0}$) and that of the synthetic control over time, $\sum_{i=}^{N}w_{i}(\mathbf{V})Y_{it}$.
Formally, $\mathbf{V}^{*}$ solves
$\min_{\mathbf{V}\in\mathcal{V}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{Nt}-w_{1}(\bm{V})Y_{1t}-\cdots-w_{N-1}(\bm{V})Y_{N-1t}\right)^{2},$ (7)
for some set $\mathcal{T_{0}}\subseteq\left\{1,2,\ldots,T_{0}\right\}$ of pre-treatment periods, or equivalently, in matrix notation,
$\min_{\mathbf{V}\in\mathcal{V}}=\left(\mathbf{Z}_{1}-\mathbf{Z}_{0}\mathbf{W^{*}(\mathbf{V})}\right)^{\prime}\left(\mathbf{Z}_{1}-\mathbf{Z}_{0}\mathbf{W^{*}}(\mathbf{V})\right)\$ (8)
where $\mathcal{V}$ is the set of all nonnegative diagonal $(K\times K)$ matrices, and $\mathbf{Z_{1}}$ and $\mathbf{Z_{0}}$ are a vector of pre-treatment outcome values for the treated unit and a matrix of pre-treatment outcomes for the donor units. Finally, the final weights for the synthetic control are given by $\mathbf{W}^{*}(\mathbf{V^{*}})$.
4.2 Inference
After estimating the effect of Explicit Instruction on NAPLAN scores in Charlestown South, we conduct a placebo test to check whether any effect we find is likely due to chance or not. To do so, we run the same SC analysis on all possible permutations of treatment assignment (keeping the timing of the treatment fixed at the original value). We run an additional 108 SC analyses, pretending in each that a different donor school was treated, rather than Charlestown South. We then use exact p-values to formally test whether the effect sizes that we find for Charlestown South are ”extreme” compared to those found in the 108 placebo analyses. If they are extreme, we take that as evidence that our results are not due to chance.
Formally, we calculate the exact p-value
$p:=\frac{\sum_{j=1}^{N}\mathbb{I}\left[RMSPE_{i}\geq RMSPE_{N}\right]}{N}$ (9)
where $\mathbb{I}[\mathbf{A}]$ is an indicator function and $RMSPE_{i}$ is the ratio mean squared prediction error, defined as
$RMSPE_{i}:=\frac{\sum_{t=T_{0}+1}^{T}\left(Y_{i,t}-\widehat{Y_{i,t}(0)}\right)^{2}/\left(T-T_{0}\right)}{\sum_{t=1}^{T_{0}}\left(Y_{i,t}-\widehat{Y_{i,t}(0)}\right)^{2}/T_{0}}$ (10)
where the numerator is the post-treatment MSPE of school $i$ and the denominator is the pre-treatment MSPE of that same school. We use the above statistic to test the sharp null
$H_{0}:Y_{i,t}(1)=Y_{i,t}(0)\text{ for each region }i\in\{1,\ldots,N\}\text{ and time period }t\in\{1,\ldots,T\}$ (11)
which will or will not be rejected based on pre-set significance level $\alpha=0.05$.

### v2

@@ Abstract
There has been substantial public debate about the potentially deleterious effects of the long-run move to “inquiry-based learning” in which students are placed at the center of an educational journey and arrive at their own understanding of what is being taught. There have been numerous calls for a return to “direct” or “explicit” instruction. This paper focuses on the impact of implementing explicit instruction via peer modelling, estimating its causal effect on student performance in standardized tests. We utilise a unique setting in Australia—a country in which all students in grades 3, 5, 7, and 9 undergo annual basic skills tests (“NAPLAN”)—and administrative data. We take a synthetic control and use Charlestown South Public School—a median-performing school—as a case study. Importantly, Charlestown South implemented explicit instruction via peer modelling, with teaching staff sitting-in during the classes of a high-performing explicit-instruction school. We find that the student performance gains are large (0.9-2.7 standard deviations) and persistent.
@@
Significance Statement
Schools across advanced economies face declining performance and debate over effective pedagogy. We study a practical, scalable implementation of Explicit Instruction: teachers from a median-performing public school in New South Wales learned the approach by observing a top-performing school and then applied it school-wide. Using administrative test data and a synthetic control design, we estimate large and persistent gains in reading and numeracy for Years 3 and 5. These results provide causal evidence that a high-fidelity, peer-modelling implementation of Explicit Instruction can substantially raise achievement at scale in a public-school setting, informing policy makers seeking cost-effective ways to improve student performance.
@@ 1 Introduction
While the source of this decline is hard to pin down, many have speculated that a long-run trend in the way students are taught is to blame. This trend has involved a move away from teachers directly imparting learning to students, toward so-called “inquiry-based learning” in which students are placed at the center of an “educational journey” where they arrive at their own understanding of what is being taught. This approach emerged from the “constructivist” movement in education which emphasizes hands-on learning aimed at solving “real-world” problems. This has led to numerous calls for a return to “direct” or “explicit” instruction (1). This is sometimes called a “back-to-basics” approach to education.
@@
The purpose of this paper is not to litigate the cause of declining educational performance in many advanced economies. Rather, we focus on estimating the causal effect of implementing the explicit instruction pedagogy via peer-modelling on student performance in standardized tests. Here, by peer-modelling implementation, we mean the process where teachers from a given school (i) learn a pedagogical approach by sitting in the classes of teachers from a school successfully implementing such approach, and (ii) use the approach in their own school. To measure this impact, we utilize a unique setting in Australia—a country in which all students in grades 3, 5, 7, and 9 undergo annual basic skills tests (“NAPLAN”)—and apply the synthetic control method (2) to administrative data. We find strong causal evidence that a peer-modelling implementation of explicit instruction improves student performance and the magnitude is substantial, with point estimates ranging between 0.9 and 2.7 standard deviations. We emphasise the implementation aspect and build the counterfactual outcomes using a pool of schools that includes schools already attempting to introduce explicit instruction in their pedagogical approach.
@@
There is a large and multi-disciplinary literature providing evidence of positive sign on the effectiveness of explicit instruction and similar pedagogical approaches. While Explicit Instruction is defined in various ways throughout the literature, we follow the definition from Archer and Hughes (3)—a classic textbook on this topic. See Section 2.1 for details. The relevant Education literature include studies on reading for primary school students (4), on mathematics in Kindergarten (5), and on using explicit instruction to teach mathematics to primary (6) and secondary school students with special needs (7). For reviews focusing on the education literature, see Archer and Hughes (3), Hughes et al. (8), and Stockard et al. (9)’s comprehensive meta-analysis.
@@
Economists have also been interested in assessing the impact of explicit and structured teaching approaches on student outcomes. Difference‐in‐differences studies of the “Literacy Hour” and synthetic-phonics reforms find substantial gains in early reading outcomes (approx. 0.20.30.2-0.3 standard deviations; s.d.) and long-run benefits for low-performing students (10, 11). Using field experiments in the US, (12) provides evidence that importing the direct instruction practices of charter schools—extended instructional time, data‐driven teaching, intensive tutoring—into public schools raises mathematics achievement by up to 0.18 s.d. one year after the implementation. Schwerdt and Wuppermann (13) estimates a fixed-effects model on US data and finds that approaches where the teacher takes centre stage and presents the content in a structured format outperform problem‐solving approaches. Jacob and Lefgren (14) uses a regression discontinuity design and focusing on mathematics, and Banerjee et al. (15) and Piper et al. (16) running RCTs in India and Kenya, respectively, to study the effect of structured approaches on numeracy and literacy skills—all finding positive impacts. Other examples of studies related to structured-teaching interventions are Romero and Sandefur (17), an RCT where 93 Liberian schools outsourced were randomly-assigned to be managed by private operators—who used scripted, highly-structured curricula—and Duflo et al. (18), where where one of the treatment arms of their RCT in Ghana was based on remedial tutoring that uses structured lesson guides. They also find positive impacts.
@@
Our paper adds to this literature in several ways. First, it shows the importance of a high-fidelity implementation of Archer and Hughes (3)’s Explicit Instruction approach. In our case study, this is achieved by the teachers of the treated school emulating the teachers of a successful Explicit-Instruction school. Specifically, the principal and the teachers of the treated school visited the successful school for three days and learned how to best implement Explicit Instruction by sitting-in during their classes (called “modelling”, in the education literature). This seems not only to be an effective solution, but also a cost-effective one, compared—for instance—to hiring dedicated coaching staff for long periods of time (see Section 2.3), and particularly in high-income countries like Australia, where the cost of labour is high. The fact that this is happening outside of an experimental setting—and within a public-school context—should also reassure policy-makers that the performance gains we estimate are attainable.
@@
Second, our intervention of interest does not target a specific subject (e.g., 10, 5, 11) or student sub-group (e.g., 19, 7, 20), instead changing the school pedagogy across all teaching areas and benefiting all students. This allows our evidence to provide insight on the broad benefits of a school-wide adoption of Explicit Instruction, compared to a partial or targeted one. Studies of charter schools and their practices (e.g., 21, 12) are similar to our study in this respect, but do not focus strictly on features of Explicit Instruction.
@@
Third, we are able to follow students for 8 years after the implementation, while even most studies with longer follow-up times are limited to less than two years (e.g., 14, 15, 19). This allows us to (i) learn that transitioning from one pedagogical approach to another can take time, and (ii) that that the benefits of Explicit Instruction are persistent in the long run. Machin and McNally (10) follow students between 1996 and 1998, but focus exclusively on impacts of reading and English skills.
@@
Synthetic control method
While the Synthetic Control Method (SCM) was first introduced by Abadie and Gardeazabal (2) to study the effect of conflict on economic output, since then it has become a mainstream method in any field working on causal questions using observational data (e.g. 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35).
We compare Charlestown South to a “synthetic twin” built as a weighted average of similar NSW schools that did not adopt the peer-modelled Explicit Instruction. The weights are chosen so that, before 2014, the synthetic twin closely matches Charlestown South on test-score trajectories and key school characteristics. After 2014, any systematic divergence in outcomes between Charlestown South and its synthetic twin is interpreted as the causal effect of the implementation. We also conduct standard placebo tests by pretending that other schools were treated to assess whether effects of similar magnitude would appear by chance.
We provide further methodological information in Section 6 and in the appendix.
@@ 2.1 The Pillars of Explicit Instruction
While (3) define sixteen Explicit Instruction principles, we report the five essential components identified in Hughes et al. (8). The first essential component consists in segmenting complex skills, where complex tasks are divided into smaller simpler units. The second is to indicate important features of the content by either showing, if an action is being taught, or telling, i.e. thinking aloud, if a concept is being taught.
@@
Note that, while some components of Explicit Instruction are found in other pedagogical approaches, Explicit Instruction is distinct from similar-sounding pedagogies such as “Direct Instruction” and “Direct Explicit Instruction” (8). For instance, Direct Instruction includes scripted lessons—which are absent in Explicit Instruction. Moreover, Direct Explicit Instruction and Direct Instruction include both curricular (what to teach) and instructional (how to teach it) elements, while Explicit Instruction purely focuses on instruction (8).
@@ 2.3 Other attempts at introducing Explicit Instruction in NSW
Before Charlestown South started implementing Explicit Instruction, the New South Wales public-school system had already introduced several structured-teaching programmes that tried to introduce some principles of Explicit Instruction. We provide some examples below.
The Early Action for Success (EaFS) policy was launched in 2012 across the Kindergarten, Year 1 and Year 2 classrooms of 228 low-performing schools. It provides target schools with “instructional leaders”, specially-hired teachers whose role is to train teacher staff in implementing explicit instruction routines both targeting the whole classroom and small groups of students who are not performing well (36). The Language, Learning and Literacy programme (L3) was piloted in 2010 and scaled to more than 900 Kindergarten classrooms by 2014, providing teachers with scripted, thirty-minute sequences on guided reading and writing (37). The Focus on Reading was rolled out state-wide in 2011 and trained teachers Years 3–6 teachers turn comprehension lessons into a deliberate, step-by-step routine. (38). All these initiatives are in line with research on evidence-based best practices in teaching conducted by the NSW Department of Education’s Centre for Education Statistics and Evaluation (CESE; 39), which recommended explicit/direct teaching approaches based on a review of the literature and was the basis of the change in pedagogical approach in NSW public schools marked by NSW Government (40).
@@ 3 Data
The NSW Department of Education dataset gives us information on: each school’s geo-location (both encoded as latitutde-longitude coordinates and as postcodes), whether it is academically selective, how remote is its surrounding area, whether it includes at least one Opportunity Class (a form of “tracking”), the year in which it hired the first teacher, school-level data on student attendance, the school’s type (primary or secondary) and its sub-type (e.g. Kindergarten to Year-6), whether it is single-sex or coeducational, if it is open until late in the and if it has a healthy canteen. It also reports time-varying school-level information such as the Index of Community Socio-Educational Advantage (ICSEA)  [^3] The ICSEA is an index that summarizes information about the occupation and earnings of children’s parents within a given school, how remote the school is and the percentage of indigenous student enrolled. , the attendance rate, the share of girls, the share of indigenous students, the share of students with a language background other than English, the average teacher-student ratio.